How do I solve:-5x+y-4z=60 How do I solve: -2x+5y+3z=-25
in variables 2x+4y+3z=-12 in variables -4x-3y-8z=-39
of three 6x-3y-2z=-52 of three 6x+8y-5z=14
How do I solve:-2x+15y+z=44 How do I solve: 8x+3y+z=43
in variables 4x+3y+3z=18 in variables -3x+5y+2z=32
of three -3+6y-z=8 of three 5x-2y+5z=24

Answer by Alan3354(69443)

How do I solve:-5x+y-4z=60 How do I solve: -2x+5y+3z=-25
in variables 2x+4y+3z=-12 in variables -4x-3y-8z=-39
of three 6x-3y-2z=-52 of three 6x+8y-5z=14
-------------------------
Enter like this:
-5x+y-4z=60
2x+4y+3z=-12
6x-3y-2z=-52
====================
You can use elimination or determinants to solve for x, y & z
==============
"in variables of three" makes no sense.